### v1

# Geometric Mechanics of Thin Periodic Surfaces

Abstract
Thin surfaces are ubiquitous in nature, from leaves to cell membranes, and in technology, from paper to corrugated containers. Structural thinness imbues them with flexibility, the ability to easily bend under light loads, even as their much higher stretching stiffness can bear substantial stresses. When surfaces have periodic patterns of either smooth hills and valleys or sharp origami-like creases this can substantially modify their mechanical response. We show that for any such surface, there is a duality between the surface rotations of an isometric deformation and the in-plane stresses of a force-balanced configuration. This duality means that of the six possible combinations of global in-plane strain and out-of-plane bending, exactly three must be isometries. We show further that stressed configurations can be expressed in terms of both the applied deformation and the isometric deformation that is dual to the pattern of stress that arises. We identify constraints rooted in symplectic geometry on the three isometries that a single surface can generate. This framework sheds new light on the fundamental limits of the mechanical response of thin periodic surfaces, while also highlighting the role that continuum differential geometry plays in even sharply creased origami surfaces.
I Introduction
Thin sheets occur widely in nature, from macroscale leaves and petals to cell membranes and viral capsids. Thin sheets are also used widely in engineering, from cutting-edge origami metamaterials [1, 2, 3] to nanotechnology such as graphene kirigami [4, 5, 6] to ubiquitous corrugated structures such as cardboard and shipping containers [7, 8, 9]. In all cases, thinness means that the systems require little volume or mass while maintaining a combination of strength and flexibility that is determined largely by their geometry.
Recent work on origami and smooth surfaces has hinted at deep mathematical relationships between different modes of deformation. In particular, even though the in-plane and out-of-plane Poisson’s ratios of thin elastic slabs are necessarily equal, for an ever-widening classes of periodically patterned surfaces, such as those shown in Fig. 1, the ratios are instead exactly opposite [10, 11, 12, 13, 14]. The structure of these linear, infinitesimal properties endures even as the ratios themselves change under nonlinear deformation [15], and may inform more complex nonlinear phenomena such as buckling, wrinkling and defect formation.
In this paper, we reveal that the vertex duality between folding modes and equilibrium stresses previously explored in triangulated origami [16, 17, 18, 19, 20, 21, 22, 23] can be extended to smooth surfaces. We introduce the concept of shape-periodic deformations that combine uniform strain and bending of periodically patterned surfaces and show that any such sheet has six such deformations, exactly three of which must be isometries. No matter how a sheet is patterned, the set of isometries is not arbitrary and instead corresponds to a Lagrangian subspace [24, 25, 26], meaning that the form of each isometry limits which others are possible.
The rest of the paper is structured as follows. In Section II we review the notions of equilibrium stresses and modes of linear isometry. We then present the duality between them, which holds for smooth, simply-connected surfaces that can deform isometrically. In Section III, we apply this duality to doubly periodic surfaces to show that each such surface has three shape-periodic isometries and that these isometries form a Lagrangian subspace. In Section IV our results are extended to creased surfaces such as origami and verified numerically. Finally, in Section V we relate our results to the literature and suggest directions for future research investigations.
II Thin sheets’ isometries, equilibrium stresses and their duality
In this work, following convention in the literature [27], we work with the two-dimensional elasticity of the shell’s mid-surface, assuming no strain along the thickness direction. Consequently, going forward, we disregard the shell’s thickness direction, and all physical quantities introduced below are defined with respect to the mid-surface—for example, stress components have units of force per unit length.
Figure 1: We consider thin sheets that are either flat (a), singly periodic (b), doubly periodic (c). In addition to such smooth sheets, our results apply to creased sheets such as origami (d). Prototypical doubly periodic sheets are readily realized through common fabrication techniques (e) and (f).
As shown in Fig. 2 (a), we consider a thin shell whose mid-surface is described by $\mathbf{x}\left(u^{1},u^{2}\right)$, with the arguments representing local coordinates that have the dimension of length. The surface, which is embedded in the conventional three-dimensional Euclidean space, in general has a nonzero local Gaussian curvature and cannot be developed from a flat sheet. We assume the surface to be smoothly differentiable almost everywhere, though our results also apply to sheets like origami, which feature sharp creases along certain lines [see Figs. 1 (d) and (f)].
II.1 Linear isometries
Isometric deformations do not locally stretch a material. For thin shells, they often correspond to low-energy deformation modes, as stretching is energetically costlier than bending [28, 29]. The relative importance of stretching compared to bending is captured by the two-dimensional Föppl-von Kármán number associated with the shell’s mid-surface [30]:
$\mathrm{FvK}\equiv\frac{\mathrm{Stretching\ stiffness}\times L_{\mathrm{defo}}^{2}}{\mathrm{Bending\ stiffness}}\approx 10\left(\frac{L_{\mathrm{defo}}}{t}\right)^{2},$ (2.1)
which is much greater than unity for thin surfaces undergoing large-scale deformations compared to their thickness (i.e., $L_{\mathrm{defo}}\gg t$). Such an isometric deformation field, denoted $\delta_{\mathrm{iso}}{\mathbf{x}}\left(u^{1},u^{2}\right)$, thus consists of local infinitesimal rotations of the reference state, so that:
$\partial_{\alpha}{\delta_{\mathrm{iso}}{\mathbf{x}}}=\bm{\omega}\times\partial_{\alpha}{\mathbf{x}},$ (2.2)
where $\bm{\omega}$ is a dimensionless angular velocity field [31] (the nomenclature does not imply dynamics), as shown in Fig. 2 (b). In fact, the compatibility of an isometry can depend only on gradients of this angular velocity, and such gradients are compatible if and only if evolving the isometry over a closed curve $\gamma$ on the surface does not change the position or orientation:
$\oint_{\gamma}\mathrm{d}{\left(\delta_{\mathrm{iso}}{\mathbf{x}}\right)}=\mathbf{0}\mbox{\quad{or} \quad}\oint_{\gamma}\mathrm{d}{\bm{\omega}}=\mathbf{0}.$ (2.3)
The angular acceleration vectors, which can be expressed in terms of the tangent vectors and the normal vectors as $\partial_{\alpha}{\bm{\omega}}\equiv{a_{\alpha}}^{\beta}\,\partial_{\beta}{\mathbf{x}}+{a_{\alpha}}\,\hat{\mathbf{n}}$, describe the variation of the angular velocity along the surface. As shown in SI Sec. II A, the closure conditions [Eq. (2.3)] imply:
$a_{\alpha}=0,$ (2.4, a)
${a_{\alpha}}^{\alpha}=0,$ (2.4, b)
$\partial_{\alpha}\left(\sqrt{g}\,\mathcal{E}^{\alpha\gamma}\,{a_{\gamma}}^{\beta}\,\partial_{\beta}{\mathbf{x}}\right)=\mathbf{0},$ (2.4, c)
where $\mathcal{E}^{\alpha\beta}\equiv\epsilon^{\alpha\beta}/\sqrt{g}$ denotes the contravariant Levi-Civita tensor, with $\epsilon^{\alpha\beta}$ the Levi-Civita symbol (whose components are $\epsilon^{12}=-\epsilon^{21}=1$ and $\epsilon^{11}=\epsilon^{22}=0$) and $g\equiv\det\left(g_{\alpha\beta}\right)$ the metric determinant [32]. As a consequence of the traceless condition Eq. (2.4, b), $\mathcal{E}^{\alpha\gamma}\,{a_{\gamma}}^{\beta}$ must be symmetric. Because of its relation to (spatial) derivatives of the angular velocity vector, we refer to ${a_{\alpha}}^{\beta}$ as the traceless angular acceleration tensor. Its use automatically incorporates the translational and rotational invariance of the metric while introducing the above nontrivial compatibility requirements. As derived in Appendix B 3 of the SI, it is closely related to the changes in the second fundamental form induced by the isometry.
Figure 2: (a) Any local portion of a smooth simply-connected surface can be represented by a vector field $\mathbf{x}\left(u^{1},u^{2}\right)$, where $u^{1}$ and $u^{2}$ are the local coordinates. The grid lines depicted in the figure are the corresponding coordinate lines, i.e., curves of constant $u^{1}$ or constant $u^{2}$. For every point on the surface there is a local frame composed of three vectors, including the unit surface normal vector ($\hat{\mathbf{n}}$) and the vectors tangent to the two coordinate lines, $\partial_{1}{\mathbf{x}}$ and $\partial_{2}{\mathbf{x}}$. In the figure we use the red-green-blue (RGB) triad of vectors to represent the local frame $\left\{\partial_{1}{\mathbf{x}},\partial_{2}{\mathbf{x}},\hat{\mathbf{n}}\right\}$ at a point. (b) When the local portion is deformed isometrically each point on the surface is displaced by a displacement vector ($\delta_{\mathrm{iso}}{\mathbf{x}}$) in such a way that infinitesimal patches of the surface are locally rotated with respect to a so-called angular velocity field ($\bm{\omega}$) without being stretched. In the figure the RGB triads of vectors still represent the local frames, but we rescale the lengths of the surface tangent vectors in the frames, so that their rescaled lengths are approximately equal to the side lengths of the corresponding infinitesimal patches. The local rotations of the infinitesimal patches can hence be understood as rotations of the corresponding local frames with respect to the angular velocity field. (c) While an isometric deformation—such as the one represented by the light-blue arrows in the figure—does not stress the surface it is mathematically dual to a stress field, as described in Section II.3. The corresponding lines of principal stress are depicted in the figure as the green and the red curves, respectively, indicating the directions along which the surface is most (the red curves) and least (the green curves) stressed. The thickness of each curve is proportional to the corresponding principal stress.
II.2 Equilibrium stresses
These conditions on the angular acceleration tensor resemble the equilibrium conditions for thin sheets. The locally generated internal forces within a thin sheet can be described in terms of a symmetric stress tensor $\sigma^{\alpha\beta}$ [28, 29]. In the case of thin sheets (membranes) whose forces are exerted in-plane (i.e., with negligible bending moments), the equilibrium conditions are given by (see SI Sec. II B):
$\partial_{\alpha}\left(\sqrt{g}\,\sigma^{\alpha\beta}\,\partial_{\beta}{\mathbf{x}}\right)=\mathbf{0}.$ (2.5)
II.3 The duality between isometries and equilibrium stresses
By comparing Eqs. (II.1) and (2.5), we arrive at the duality, which is a key result of this paper. For any given infinitesimal isometry with specified angular acceleration tensor, an equilibrium stress may be generated via:
$\sigma^{\alpha\beta}=\left(\mathrm{const.}\right)\mathcal{E}^{\alpha\gamma}\,{a_{\gamma}}^{\beta},$ (2.6)
where the constant must have units of force. The equation is easily inverted, so that an equilibrium stress corresponds to a unique isometry. Note that although each isometry can be mathematically mapped onto an equilibrium stress, deforming a sheet isometrically generates no stress. This correspondence mirrors that of the vertex duality of triangulated origami [16, 17, 18, 19, 20, 21, 22, 23].
This duality is depicted in Fig. 2 (c). The red and green lines in the figure respectively depict the local maximum and minimum stresses of a force equilibrium state of the surface, with line thickness proportional to the stress magnitudes. According to the isometry-stress duality [Eq. (2.6)], the corresponding equilibrium stress can be mapped onto an angular acceleration tensor, which can then be integrated once over the surface to obtain the angular velocity field, and integrated a second time to yield the isometric displacement field $\delta_{\mathrm{iso}}{\mathbf{x}}$, shown as the blue arrows. In this way, once an equilibrium stress is known, an isometry can be generated, and vice versa. This stress/isometry duality appears analogous to those developed by Calladine for thin shells [33] and subsequently extended to more general surfaces (see, e.g., Ref. 34).
III The relation between rigidity and flexibility of smooth periodic surfaces
Figure 3: Geometry of doubly periodic surfaces and their shape-periodic deformations. (a) A doubly periodic surface is generated by repeatedly translating the unit cell (denoted by $\mathrm{C}$), which is enclosed by the magenta curve $\partial{\mathrm{C}}$ in the figure, along the directions of $\bm{\ell}_{1}$ and $\bm{\ell}_{2}$. The coarse-grained geometry of the doubly periodic surface is flat. The corresponding flat plane is spanned by $\bm{\ell}_{1}$ and $\bm{\ell}_{2}$ and has the normal vector $\hat{\mathbf{z}}$. We choose as the zero-height level the mid-height plane of the periodic surface, which coincides with the coarse-grained flat plane. (b) Under a shape-periodic deformation, the Euclidean distance between any pair of points on the deformed surface remains invariant if both points are shifted by the same number of unit cells; e.g., $\left\lVert{\mathbf{M}-\mathbf{N}}\right\rVert=\left\lVert{\mathbf{M}^{\prime}-\mathbf{N}^{\prime}}\right\rVert$ in the figure. The inset illustrates the corresponding deformation of the coarse-grained flat plane. (c) The deformation of the coarse-grained flat plane can be decomposed into six deformation modes: three in-plane modes and three out-of-plane modes, as illustrated by $\mathrm{(c,i)}$ and $\mathrm{(c,ii)}$, respectively.
III.1 Smooth periodic surfaces
This duality has important implications for the response of surfaces that are doubly periodic; i.e., those invariant under two linearly independent translations. Let $\bm{\ell}_{1(2)}\equiv\ell^{1(2)}\,\hat{\mathbf{\bm{\ell}}}_{1(2)}$ denote the two translation vectors. We can thus always generate a doubly periodic surface by first choosing the geometry of its unit cell (denoted by $\mathrm{C}$) and then repeatedly translating the unit cell by $\bm{\ell}_{1}$ and $\bm{\ell}_{2}$. [See Figs. 1 and 3 (a)]. The geometry of the unit cell can be characterized by some doubly periodic vector-valued function $\mathbf{x}^{\mathrm{p}}\left(u^{1},u^{2}\right)$, satisfying the periodic conditions:
$\mathbf{x}^{\mathrm{p}}\left(u^{1},u^{2}\right)$ $=\mathbf{x}^{\mathrm{p}}\left(u^{1}+\ell^{1},u^{2}\right),$ (3.1, a)
$\mathbf{x}^{\mathrm{p}}\left(u^{1},u^{2}\right)$ $=\mathbf{x}^{\mathrm{p}}\left(u^{1},u^{2}+\ell^{2}\right).$ (3.1, b)
And the corresponding doubly periodic surface can be accordingly parameterized as:
$\mathbf{x}\left(u^{1},u^{2}\right)\equiv u^{\alpha}\,\hat{\mathbf{\bm{\ell}}}_{\alpha}+\mathbf{x}^{\mathrm{p}}\left(u^{1},u^{2}\right).$ (3.2)
In the expression above, the surface can be thought of as a (macroscopic) smooth plane and a (microscopic) periodic unit cell, the latter of which we take to average to zero. Consequently, since the surface macroscopically resembles a plate—specifically, the one spanned by its translation vectors—the macroscopic deformations can be described using the concepts of plate deformations. The following notations are used to characterize the geometry of the macroscopic plate corresponding to a periodic surface: $\phi$ denotes the angle between the translation vectors; $\hat{g}_{\alpha\beta}\equiv\hat{\mathbf{\bm{\ell}}}_{\alpha}\cdot\hat{\mathbf{\bm{\ell}}}_{\beta}$ denotes the planar metric components with determinant $\hat{g}\equiv\det\left(\hat{g}_{\alpha\beta}\right)=\sin^{2}{\phi}$; $\hat{\mathcal{E}}^{\alpha\beta}\equiv\epsilon^{\alpha\beta}/\sqrt{\hat{g}}$ denotes the planar contravariant Levi-Civita tensor; and $\hat{\mathbf{z}}\equiv\hat{\mathbf{\bm{\ell}}}_{1}\times\hat{\mathbf{\bm{\ell}}}_{2}/\sqrt{\hat{g}}$ denotes the unit normal vector of the plate.
III.2 Shape-Periodic deformations
We focus on deformations that induce the same shape in every unit cell, which we thus refer to as shape-periodic. In a shape-periodic deformed state of a periodic surface, the Euclidean distance between any pair of points is invariant if we shift both of them by an arbitrary number of unit cells, just like in the undeformed state. [See Fig. 3 (b).] Mathematically, let $(s,t)$ and $(S,T)$ denote the local coordinates of a pair of points on the periodic surface $\mathbf{x}$, and we write a deformed state of $\mathbf{x}$ as ${\mathbf{x}^{\prime}}\left(u^{1},u^{2}\right)\equiv{\mathbf{x}}\left(u^{1},u^{2}\right)+\delta{\mathbf{x}}\left(u^{1},u^{2}\right)$. By our definition, a shape-periodic deformation thus has to satisfy:
$\begin{split}&\left\lVert{\mathbf{x}^{\prime}\left(S,T\right)-\mathbf{x}^{\prime}\left(s,t\right)}\right\rVert\\[2.5pt] &\quad=\left\lVert{\mathbf{x}^{\prime}\left(S+n_{1}\ell^{1},T+n_{2}\ell^{2}\right)-\mathbf{x}^{\prime}\left(s+n_{1}\ell^{1},t+n_{2}\ell^{2}\right)}\right\rVert\end{split}$ (3.3)
for any integer tuple $(n_{1},n_{2})$. Rigid-body displacements (translations and rotations) are trivially shape-periodic but do not deform the surface.
In simple terms, a shape-periodic deformation causes a periodic surface to deform uniformly, with each unit cell of the surface experiencing the same deformation. Based on the notion of separation of length scales, such a deformation can therefore be microscopically characterized by a doubly periodic displacement field $\delta_{\mathrm{SP}}{\mathbf{x}}^{\mathrm{p}}\left(u^{1},u^{2}\right)$, which assumes the same value in every unit cell. The macroscopic part of the shape-periodic deformation corresponds to the uniform deformation modes of the macroscopic plate associated with the periodic surface. According to plate theory, these uniform deformations typically consist of uniform in-plane stretching and shearing, as well as out-of-plane bending and twisting. Consequently, a shape-periodic deformation can be macroscopically described using two constant symmetric matrices, $\left(E^{\alpha\beta}\right)$ and $\left(H^{\alpha\beta}\right)$, which quantify the constant in-plane strains and out-of-plane curvature changes of the macroscopic plate, respectively. [See Fig. 3 (c).]
Let $x_{\alpha}\equiv\mathbf{x}\cdot\hat{\mathbf{\bm{\ell}}}_{\alpha}$ and $x_{z}\equiv\mathbf{x}\cdot\hat{\mathbf{z}}$ denote the macroscopic in-plane and out-of-plane components of the periodic surface $\mathbf{x}$, respectively. Based on the expressions for the uniform deformations of a plate [29, 28], a generic shape-periodic displacement field may therefore be expressed as:
$\begin{split}\delta_{\mathrm{SP}}{\mathbf{x}}\left(u^{1},u^{2}\right)&\equiv E^{\alpha\beta}x_{\alpha}\,\hat{\mathbf{\bm{\ell}}}_{\beta}\\[2.5pt] &\quad+\frac{1}{2}\,H^{\alpha\beta}x_{\alpha}\,x_{\beta}\,\hat{\mathbf{z}}-H^{\alpha\beta}x_{\alpha}\,x_{z}\,\hat{\mathbf{\bm{\ell}}}_{\beta}\\[2.5pt] &\quad\quad+\delta_{\mathrm{SP}}{\mathbf{x}}^{\mathrm{p}}\left(u^{1},u^{2}\right).\end{split}$ (3.4)
$E$ and $H$ are the changes to the macroscopic (i.e., averaged over the unit cells) first and second fundamental forms, respectively. In SI Sec. III B 4, we further discuss the physical meaning of each term in Eq. (3.4) and explicitly show that the equation satisfies the formal definition of shape-periodic deformations [Eq. (3.3)] to linear order.
If a shape-periodic deformation is also microscopically isometric, the concept of an angular velocity field, as discussed in Section II.1, can be used to describe it, alongside Eq. (3.4). Due to the separation of length scales, the angular velocity field corresponding to an isometric shape-periodic deformation, $\bm{\omega}_{\mathrm{SP}}\left(u^{1},u^{2}\right)$, naturally consists of both microscopic and macroscopic parts. As with the surface embedding itself [see Eq. (3.2)], the microscopic part, describing the identical local rotations of the infinitesimal area elements within each unit cell, can be captured by a doubly periodic function $\bm{\omega}_{\mathrm{SP}}^{\mathrm{p}}\left(u^{1},u^{2}\right)$. The macroscopic part of $\bm{\omega}_{\mathrm{SP}}$ characterizes the relative rotations between each unit cell and its neighboring unit cells. By shape periodicity, these relative rotations must also be identical for every unit cell and can thus be captured by a pair of constant angular acceleration vectors, $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$, which characterize the constant variations across unit cells along the two translation directions. Based on these facts, the angular velocity field associated with an isometric shape-periodic deformation can be parameterized as follows:
$\bm{\omega}_{\mathrm{SP}}\left(u^{1},u^{2}\right)\equiv u^{\alpha}\,\mathbf{w}_{\alpha}+\bm{\omega}_{\mathrm{SP}}^{\mathrm{p}}\left(u^{1},u^{2}\right).$ (3.5)
This form also follows from the assumption of a periodic angular velocity tensor. In SI Sec. IV B 1, we show that the constant angular acceleration vectors $\mathbf{w}_{\alpha}$ and the curvature changes $H^{\alpha\beta}$, both describing the macroscopic out-of-plane isometric deformations of a periodic surface, are related as follows:
${\mathbf{w}}_{\alpha}=\hat{\mathcal{E}}^{\beta\gamma}\,{H}_{\gamma\alpha}\,\hat{\mathbf{\bm{\ell}}}_{\beta},$ (3.6)
where $H_{\gamma\alpha}=\hat{g}_{\gamma\mu}\,\hat{g}_{\alpha\nu}\,H^{\mu\nu}$, with $\hat{g}_{\alpha\beta}$ denoting the planar metric components.
To summarize, any shape-periodic deformation mode can be characterized by the six macroscopic quantities $E_{\alpha\beta}$ and $H_{\alpha\beta}$, together with a microscopic periodic function $\delta{\mathbf{x}_{\mathrm{SP}}^{\mathrm{p}}}$. It is useful to represent the former as a single Voigt-type [35] vector $\mathbf{v}$, whose exact form will become convenient later:
$\mathbf{v}^{\intercal}\equiv\left(\frac{1}{\sqrt{\hat{g}}}H_{22},\frac{1}{\sqrt{\hat{g}}}H_{11},-\frac{1}{\sqrt{\hat{g}}}H_{12},E_{11},E_{22},{2}E_{12}\right),$ (3.7)
where $\hat{g}$ denotes the determinant of the planar metric tensor. The vector space formed by all such six-dimensional vectors is called the deformation phase space.
[…]
The existence of exactly three linear isometries is surprising. For example, the flat surface [Fig. 1 (a)] has an infinite number of linear isometries corresponding to small local displacements in the transverse direction. However, any such purely local deformation induces no global strain or curvature. Instead, the three global isometries of the surface consist of the three independent components of $H^{\alpha\beta}$, with the sheet curling up like a cylinder or a dome.
Thus, contrary to the common perception that adding corrugation to cardboard or curving a pizza slice so it does not bend downward adds strength to the structure, we find that any (periodic) sheet will always have three global, shape-periodic isometries and three shape-periodic deformations that are not isometries. We refer to this result as conservation of flexibility.
III.3 Deformation energy and boundary work
Having characterized the deformation modes of interest, we now investigate how they couple, leading to the relation between the rigidity and flexibility of a periodic surface. We begin by formulating the energies incurred by an energy-minimizing deformation.
Recall that we consider small elastic deformations in the limit of large Föppl-von Kármán number [Eq. (2.1)], where stretching dominates over bending. In this regime, the corresponding energy is typically captured by the following quadratic membrane energy functional [34]:
$E_{\mathrm{m}}=\frac{1}{2}\,\iint_{\mathrm{C}}\mathrm{d}{\mathcal{A}}\ C^{\alpha\beta\gamma\rho}\,\varepsilon_{\gamma\rho}\,\varepsilon_{\alpha\beta}=\frac{1}{2}\,\iint_{\mathrm{C}}\mathrm{d}{\mathcal{A}}\ \sigma^{\alpha\beta}\,\varepsilon_{\alpha\beta}.$ (3.8)
Here, the integral is taken over a unit cell ($\mathrm{C}$) of a periodic surface, and $\mathrm{d}{\mathcal{A}}$ representing its area element. The symbol $\varepsilon_{\alpha\beta}$ denotes the strain tensor, which quantifies the change in the surface metric induced by the deformation. In the elastic regime, the strain tensor is linearly related to the stress tensor via the generalized Hooke’s law: $\sigma^{\alpha\beta}=C^{\alpha\beta\gamma\rho}\,\varepsilon_{\gamma\rho}$, where $C^{\alpha\beta\gamma\rho}$ denotes the components of the rank-four two-dimensional (proportional to the thickness of the sheet) stiffness tensor, which encodes the elastic properties of the surface material, such as Young’s modulus and Poisson’s ratio.
At equilibrium, the internal stresses in the surface bulk balance out, satisfying the equilibrium equations [Eq. (2.5)]. As a result, by the principle of energy conservation, the deformation energy stored in the bulk must equal the net work done by the deformation on the boundary. As shown in SI Sec. II B, the stored energy can be expressed as:
$E_{\mathrm{m}}^{\mathrm{eq}}=W=\frac{1}{2}\oint_{\partial\mathrm{C}}\mathrm{d}{u^{\gamma}}\,\left\langle{\mathcal{E}_{\alpha\gamma}\,\sigma^{\alpha\beta}\,\partial_{\beta}{\mathbf{x}},\delta{\mathbf{x}}}\right\rangle,$ (3.9)
where $\partial{\mathrm{C}}$ denotes the boundary of the unit cell $\mathrm{C}$, $\mathcal{E}_{\alpha\gamma}$ is the inverse of the contravariant Levi-Civita tensor, and $\delta{\mathbf{x}}$ is the displacement field associated with the deformation.
The stress tensor $\sigma^{\alpha\beta}$ in Eq. (3.9) corresponds to an energy-minimizing deformation. For a periodic surface, the isometry-stress duality [Eq. (2.6)] enables us to express the stored energy in terms of the angular velocity field associated with an isometric shape-periodic mode of the surface, as follows:
[…]
where the tilde notation is henceforth used to denote quantities associated with the virtual isometric mode mapped from an energy-minimizing deformation. This represents the actual physical energy as a purely geometric quantity, given in terms of an isometric deformation.
As a special case of Eq. (3.10), when there is a prestress and the imposed deformation is itself isometric—thus incurring no energy cost—the equation implies (see SI Sec. IV C 2 for further discussion):
$\oint_{\partial\mathrm{C}}\mathrm{d}{u^{\alpha}}\left\langle{\partial_{\alpha}{\bm{\omega}_{\mathrm{SP}}^{a}},\delta_{\mathrm{iso}}{\mathbf{x}_{\mathrm{SP}}^{b}}}\right\rangle=0,$ (3.11)
where the indices $a$ and $b$ label two arbitrary isometric shape-periodic modes of the periodic surface.
III.4 The relation between rigidity and flexibility
The coupling between an isometric mode and its corresponding energy-minimizing mode, as well as with a second isometric mode, is described by Eqs. (3.10) and (3.11), respectively.
[…]
where, intriguingly, all the microscopic periodic components are integrated out. The fact that the couplings depend solely on the macroscopic quantities implies that all doubly periodic surfaces can be treated macroscopically as plates, with their mechanical properties modified by the geometries of their unit cells—a notion also established using homogenization techniques (see, e.g., Refs. 36 and 37).
Of exceptional interest is the case, mentioned previously, in which the deformation is itself isometric. In that case, the energy vanishes and, upon dividing out constant terms, we have one of our main results, which we term the surface mode compatibility condition. For any two isometric modes of a single periodic surface $a,b$:
$\epsilon^{\alpha\mu}\,\epsilon^{\beta\nu}\left(H^{a}_{\mu\nu}\,E^{b}_{\alpha\beta}-H^{b}_{\mu\nu}\,E^{a}_{\alpha\beta}\right)=0.$ (3.13)
In the vector representation introduced previously [Eq. (3.7)], the couplings assume the following form:
[…]
$0$ $={\mathbf{v}}_{a}^{\intercal}\,\mathbf{J}\,\mathbf{v}_{b}^{\phantom{\intercal}},$ (3.15)
where $\mathbf{J}$ denotes the canonical symplectic matrix [38]:
$\mathbf{J}\equiv\begin{pmatrix}\phantom{-}\mathbf{0}_{3\times 3}&\mathbf{1}_{3\times 3}\\[2.5pt] {-}\mathbf{1}_{3\times 3}&\mathbf{0}_{3\times 3}\end{pmatrix}$. (3.16)
The fact that the couplings between shape-periodic modes can be expressed as symplectic inner products, as shown in Eqs. (3.14) and (3.15), means that the deformation phase space possesses a natural symplectic structure, akin to the phase space of classical Hamiltonian mechanics. Specifically, the curious combination $\left(H_{22},H_{11},-H_{12}\right)/\sqrt{\hat{g}}$ in Eq. (3.7)—more compactly expressed as $\Sigma^{\alpha\beta}\equiv\hat{\mathcal{E}}^{\alpha\mu}\,\hat{\mathcal{E}}^{\beta\nu}\,H_{\mu\nu}$—plays the role of generalized coordinates in classical mechanics. The macroscopic strain components $E_{\alpha\beta}$, with lowered indices, are conjugate to $\Sigma^{\alpha\beta}$, thereby serving as the conjugate momenta associated with the generalized coordinates.
In the symplectic language, as shown in Eq. (3.14), the energy cost incurred by an energy-minimizing shape-periodic mode can be interpreted geometrically as the phase-space area of the parallelogram formed by its associated vector $\mathbf{v}$ and that of the corresponding isometric mode, $\widetilde{\mathbf{v}}$. More remarkably, the mode compatibility condition Eq. (3.15) implies that the subspace formed by the three isometric shape-periodic modes is Lagrangian, with the property that the symplectic inner product between any pair of its elements vanishes [24, 25, 26].
The mode compatibility condition constrains the possible forms of isometries for a given surface, thereby restricting its intrinsic responses to stretching and bending—particularly when some of its isometry data are known. Specifically, if a given vector $\mathbf{v}$ corresponds to an isometric mode, then, based on Eq. (3.15), the mode associated with the vector $\mathbf{J}\,\mathbf{v}$ will inevitably incur an energy cost.
For concreteness, consider singly corrugated surfaces as an example [see Fig. 1 (b)]. Uniformly stretching such surfaces along the corrugation direction (and flattening the corrugations) does not incur an energy cost and is therefore an isometric mode. By aligning the first translation vector $\bm{\ell}_{1}$ with the corrugation direction, we can represent this isometry as $\mathbf{v}_{\mathrm{s}}^{\intercal}\equiv\left(0,0,0,E_{11},0,0\right)$. Given this information, Eq. (3.15) implies that bending the surface along the second translation vector $\bm{\ell}_{2}$, represented by $\mathbf{v}_{\mathrm{b}}^{\intercal}\equiv\left(H_{22},0,0,0,0,0\right)$, cannot be isometric, as consistent with everyday experience (e.g., bending a corrugated piece of paper). More generally, Eq. (3.15) goes beyond this simple scenario, showing that no periodic sheet—regardless of its unit cell geometry—can simultaneously exhibit the uniaxial strain mode of the corrugated sheet in this example and the transverse uniaxial bending mode of a flat sheet.
In the special case where the isometric modes of a periodic surface involve only pure stretching or pure bending, the mode compatibility condition [Eq. (3.13)] reduces to:
$\begin{split}0=\epsilon^{\alpha\mu}\,\epsilon^{\beta\nu}\,E^{a}_{\alpha\beta}\,H^{b}_{\mu\nu}=E^{a}_{11}\,H^{b}_{22}+E^{a}_{22}\,H^{b}_{11}-2\,E^{a}_{12}\,H^{b}_{12},\end{split}$ (3.17)
which recovers the result reported in Ref. 39, obtained via homogenization techniques and the method of averaging [40]. Consequently, as established in Ref. 39, a coordinate transformation that sets either $E^{a}_{12}$ or $H^{b}_{12}$ to zero causes the mode compatibility condition to imply that the surface possesses opposite macroscopic in-plane and out-of-plane Poisson’s ratios:
$\hat{\upsilon}_{\mathrm{in}}\equiv\frac{E^{a}_{2^{\prime}2^{\prime}}}{E^{a}_{1^{\prime}1^{\prime}}}=-\frac{H^{b}_{2^{\prime}2^{\prime}}}{H^{b}_{1^{\prime}1^{\prime}}}\equiv-\hat{\upsilon}_{\mathrm{out}},$ (3.18)
where the primed indices refer to the transformed coordinates.
IV Origami and triangulation of smooth surfaces
Figure 4: Numerically computed isometries of surfaces. (a) The surfaces and their triangulations used in the simulations: a Miura-ori (first row), a graph of translation (second row) and a perturbed graph of translation (third row). Explicit parametrizations of the latter two surfaces are provided in Fig. 2 of the SI. Unit cell geometries are shown in the corresponding insets. The unit cell of the triangulated Miura-ori comprises eight triangular faces, whereas each of the other two triangulated surfaces has $2\mathcal{N}^{2}=288$ triangular faces in its unit cell, where $\mathcal{N}$ denotes the number of grid squares per row (and per column) and serves as the parameter controlling the triangulation density. (b) The three numerically computed linear isometries are shown in the three columns, along with their vector representations [of the form in Eq. (3.7)] and the corresponding macroscopic plate deformations, displayed in the insets.
Although our analysis has assumed smooth surfaces, the analytical results for periodic surfaces—Eqs. (3.12) and (3.13)—also apply to piecewise smooth geometries, such as origami tessellations and triangulated surfaces that are smooth except at sharply creased interfaces [Fig. 4 (a)]. As discussed in SI Sec. VII A, this result can be derived by decomposing the line integral around the unit cell boundary into integrals along the creases, rendering the relevant quantities smooth within the regions bounded by the creases, following the approach of Ref. 39.
Origami is a rich and vibrant field [41, 42], and triangulations are commonly used to model the isometries of origami tessellations, including those with non-triangular faces, to capture face bending [43, 44, 45, 16]. Origami and triangulated surfaces thus serve as a natural testbed for our results: the presence of three isometric modes that obey the mode compatibility condition [Eqs. (3.13) or (3.15)]. As detailed in SI Sec. VII B 2, the isometries of a triangulated surface can be characterized by changes in the dihedral angles between adjoining triangular faces, subject to the requirement that the total change around any closed loop satisfies the closure conditions in Eq. (2.3). In this formulation, the scalar dihedral angle changes serve the role of the angular acceleration tensor introduced around Eqs. (II.1).
As shown in the first row of Fig. 4, a triangulated Miura-ori indeed generates three isometries that lead to global strains and curvatures whose forms are consistent with those reported in the literature [11, 10]. With the six-dimensional vectors [of the form given in Eq. (3.7)] corresponding to these isometries, as shown in the insets, one can verify that the mode compatibility condition is satisfied within numerical precision for each pair of isometries. This confirmation not only validates our theory but also demonstrates that the assumption that bending is concentrated at the creases is exact. In this way, we show that our results apply not only to strictly smooth surfaces but also to origami tessellations.
The second row of Fig. 4 corresponds to triangulations of a graph of translation [46], for which two of the three isometric modes are known analytically [47]. In contrast to the case of triangulated Miura-ori, we now use a finite number of triangles to approximate an ideal curved surface. In Fig. 5, we observe that, as the number of triangles used increases, the isometries always obey the mode compatibility condition and gradually converge to a fixed subspace, indicating that the isometries of the graph of translation are faithfully reproduced. Indeed, two of the three isometries that can be determined analytically lie within the subspace. Finally, the last row of Fig. 4 depicts a perturbed graph of translation whose isometries are not known analytically and involve both macroscopic stretching and bending. As illustrated in Fig. 5, our results hold for the surface. This intrinsic coupling between stretching and bending arises for generic surfaces, in contrast with the modes considered in Ref. 39, which were either pure bending or pure stretching.
While such triangulations can, in principle, approximate any smooth surface, numerical limitations may hinder the convergence of isometric subspaces for arbitrary geometries, possibly due to rigidifying curves [48] that arise near regions of vanishing Gaussian curvature. Such regions must exist in any of the periodic surfaces considered here, as the Gauss-Bonnet theorem implies that the average Gaussian curvature over a unit cell vanishes—even under isometric deformations, since Gaussian curvature is intrinsic.
Figure 5: Convergence of isometric subspaces and validation of the mode compatibility condition. The red and cyan curves correspond to the red and cyan surfaces in Fig. 4, respectively. For each triangulated surface, $\mathcal{N}$ denotes the number of grid squares per row and per column, so that the unit cell contains $2\mathcal{N}^{2}$ triangular faces, as seen in the insets. The solid curves show the values obtained by substituting the isometry data into the mode compatibility condition [Eqs. (3.13) or (3.15)], averaged over all non-redundant isometry pairs. The dashed curves depict the convergence of isometric subspaces for the triangulated surfaces. Each data point measures the “distance” between subspaces corresponding to $\mathcal{N}$ and $\mathcal{N}-1$ grid squares per row and column, quantified by $d^{\mathbb{R}^{6}}_{\mathcal{N},\,\mathcal{N}-1}$, a Euclidean-based metric for comparing equal-dimensional subspaces in $\mathbb{R}^{6}$ (see SI Sec. VII C for details and for convergence behavior using other types of metrics).
V Discussion
We have uncovered a general duality associated with any piecewise smooth thin membrane surface: Any isometric deformation can be mapped to an equilibrium stress (and thus energy-minimizing deformation), and vice-versa. For periodic surfaces, the duality establishes that, within the six-dimensional space of combined macroscopic in-plane (coarse-grained strain) and out-of-plane (coarse-grained bending) deformation modes, exactly three modes (and their linear combinations) must correspond to microscopic isometries, thereby being energetically favored. Moreover, we find that the three isometric modes couple symplectically with each other, satisfying the surface mode compatibility condition [Eqs. (3.13) or (3.15)]. In the special case for which the isometric modes consist of pure macroscopic strains or curvature changes, the mode compatibility condition reduces to the exciting result recently established in the literature [39], thus providing a mathematical foundation for the equal and opposite macroscopic in-plane and out-of-plane Poisson’s ratios possessed by several types of origami tessellations—including the Miura-ori [11, 10], the eggbox [12] and the morph pattern [13].
Our main results for periodic surfaces—namely, the symplectic energy-cost formula [Eqs. (3.12) or (3.14)] and the mode compatibility condition—are universal in the sense that neither expression depends on the microscopic details of the unit cells of a periodic surface. Despite this, it is important to emphasize that, unlike other existing theories for origami tessellations [49], the theory developed in this work is an exact microscopic theory that does not rely on any coarse-graining or homogenization techniques.
There are a few puzzles in our work that remain to be solved. First, as mentioned in Section IV, it remains unclear why the isometric Lagrangian subspace associated with a general triangulated periodic surface does not converge as the number of triangles increases, even though the mode compatibility condition is consistently obeyed. Second, the revealed symplectic structure of the deformation phase space and the Lagrangian nature of its isometric subspace point to deeper underlying principles that are not yet fully understood. It would also be interesting to investigate whether these structures, as well as the isometry-stress duality, persist when the assumptions of membrane states and small deformations are relaxed.
The theoretical framework developed in this work can be extended in the following directions. First, similar to the direction pursued for origami tessellations [16], one could incorporate the constraints imposed by geometric nonlinearities into the theory to investigate which linear isometric modes of a periodic surface extend to the nonlinear regime, where microscopic deformations are no longer small. Another promising direction would be to introduce extra microscopic structures, such as symmetries, fluctuations or disorder, into the unit cells of a periodic surface and examine their effects on the surface’s macroscopic behavior. Finally, perhaps of greater practical interest is the inverse problem: How to design unit cells that realize a desired set of isometries permitted by the mode compatibility condition.
In conclusion, our work reveals the intrinsic mechanical behaviors of general periodic surfaces, and we anticipate that our results, along with future investigations, will prove valuable for the industrial design of corrugated structures, with specific applications such as cardboard, shipping containers and other related materials.
VI ACKNOWLEDGMENTS
The authors acknowledge financial support from the Army Research Office through the MURI program (#W911NF2210219) and through the National Science Foundation CAREER program (#2338492), NSF CMMI- 2247095 (CDS) and through NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). We would also like to thank Evan Dickson for fabricating the 3D-printed surface and folding the Miura-ori shown in Fig. 1.

### v2

@@
 $\begin{split}\delta_{\mathrm{SP}}{\mathbf{x}}\left(u^{1},u^{2}\right)&\equiv E^{\alpha\beta}x_{\alpha}\,\hat{\mathbf{\bm{\ell}}}_{\beta}\\[2.5pt] &\quad+\frac{1}{2}\,H^{\alpha\beta}x_{\alpha}\,x_{\beta}\,\hat{\mathbf{z}}-H^{\alpha\beta}x_{\alpha}\,x_{z}\,\hat{\mathbf{\bm{\ell}}}_{\beta}\\[2.5pt] &\quad\quad+\delta_{\mathrm{SP}}{\mathbf{x}}^{\mathrm{p}}\left(u^{1},u^{2}\right).\end{split}$ (3.4)
-$E$ and $H$ are the changes to the macroscopic (i.e., averaged over the unit cells) first and second fundamental forms, respectively.
-In SI Sec.
+Given the form of a shape-periodic mode $\delta_{\mathrm{SP}}{\mathbf{x}}\left(u^{1},u^{2}\right)$, the corresponding changes in the macroscopic (i.e., unit-cell-averaged) first and second fundamental forms, $E_{\alpha\beta}$ and $H_{\alpha\beta}$, can be obtained using:
+$H_{\alpha\beta}$ $=\left\langle{{\mathfrak{d}}_{\alpha}{{\mathfrak{d}}_{\beta}{\delta_{\mathrm{SP}}{\mathbf{x}}}},\hat{\mathbf{z}}}\right\rangle$, (3.5, a)
+$E_{\alpha\beta}$ $=\frac{1}{2}\left(\left\langle{{\mathfrak{d}}_{\alpha}{{\delta_{\mathrm{SP}}{\mathbf{x}}}},\hat{\mathbf{\bm{\ell}}}_{\beta}}\right\rangle+\left\langle{{{\mathfrak{d}}_{\beta}{\delta_{\mathrm{SP}}{\mathbf{x}}}},\hat{\mathbf{\bm{\ell}}}_{\alpha}}\right\rangle\right)+H_{\alpha\beta}\,x_{z},$ (3.5, b)
+where the lattice derivative of a function $f\left(u^{1},u^{2}\right)$ is defined as ${\mathfrak{d}}_{1}{f}\equiv\left[{f}\left(u^{1}+\ell^{1},u^{2}\right)-{f}\left(u^{1},u^{2}\right)\right]/\ell^{1}$ (and similarly for ${\mathfrak{d}}_{2}{f}$).
+We note that Eqs.
+(III.2) exactly mirror the usual expressions used to compute the microscopic infinitesimal strains and curvature changes.
+[See SI Sec.
+VI B for a discussion of the gauge-invariance properties of the macroscopic strains $E_{\alpha\beta}$, which give rise to the term $H_{\alpha\beta}\,x_{z}$ in Eq.
+(3.5, b).] In SI Sec.
 III B 4, we further discuss the physical meaning of each term in Eq.
@@
 (3.8)] and the corresponding macroscopic plate deformations, displayed in the insets.
+The macroscopic strains and curvature changes associated with each mode are computed using Eqs.
+(III.2), with the surface average heights set to zero, as shown for example in Fig.
+3 (a).
+(See SI Sec.
+VII B 3 for more details.)
 Although our analysis has assumed smooth surfaces, the analytical results for periodic surfaces—Eqs.
@@
 (II.1).
+In our simulations, after triangulating the selected surfaces, we obtain their isometry data—the dihedral angle changes and, subsequently, the isometric displacements—by computing the zero modes of their associated compatibility matrices (see SI Sec.
+VII B 2 for details).
+With the isometry data in hand, Eqs.
+(III.2) are used to extract the macroscopic strains and curvature changes corresponding to the isometries.
+A caveat applies when we report the numbers: Because the macroscopic strains are gauge-dependent (see SI Sec.
+VI B for details), the average heights of the selected surfaces are set to zero to produce physically intuitive results; in particular, in this height gauge, bending a surface leaves its mid-height plane strainless.
 As shown in the first row of Fig.
@@
 VI ACKNOWLEDGMENTS
-The authors acknowledge financial support from the Army Research Office through the MURI program (#W911NF2210219) and through the National Science Foundation CAREER program (#2338492), NSF CMMI- 2247095 (CDS) and through NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).
+The authors acknowledge financial support from the Army Research Office through the MURI program #W911NF2210219 (DZR) and through the National Science Foundation CAREER program #2338492 (DZR), NSF CMMI- 2247095 (CDS) and through NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP) (DZR).
 We would also like to thank Evan Dickson for fabricating the 3D-printed surface and folding the Miura-ori shown in Fig.